PLEASE ANSWER FAST !!!!!
 
Two  identical spheres A and B each having a charge of 6.5 × 10-7 C are separated by a distance of 50 cm in air. A third uncharged sphere of same size is brought in contact with the first then brougth in contact with the second and afterwards removed from both. What is the new electrical force between the first two spheres . 

Grade:12

1 Answers

Arun
6 years ago

Charge on sphere A, qA = Charge on sphere B,

qB = 6.5 × 10−7 C

Distance between the spheres, r = 50 cm = 0.5 m

Force of repulsion between the two spheres,

F =k* qa *qb / r^2

Where, ∈0 = Free space permittivity

1/ 4* pi* epsilon=9x109N mC2

∴F = 9 * 10^9 * (6.5* 10^ –7)^2 / 0.25

=1.52X10-2N Therefore, the force between the two spheres is 1.52 × 10−2 N.